Question 8 Given a2 = 6 and an = 2an-1 + 3n find the third and fourth term.

PLEASE SHOW WORK
Answer #1

solution a = 6 an=2anul +3n Third Term an=3 Az = 2 a 3-1 +3(3) Az = 2q +9 az = 2(6) +9 az = 12+9 az - 21 Fourth Term ין nay a